the equation of a parabola with the given focus and directrix
F(0,-5), y=5

Answer by lwsshak3(6481)

the equation of a parabola with the given focus and directrix
F(0,-5), y=5
This is an equation of a parabola that opens downwards. (directrix at y=5, shows this)
Its standard form: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of vertex.
For given equation:
axis of symmetry: x=0
x-coordinate of vertex=0 (from focus)
y-coordinate of vertex=0 (half way between -5 and 5 on the axis of symmetry
p=5 (distance to vertex from focus or directrix on the axis of symmetry
4p=20
equation of given parabola:
x^2=20y